OpenStudy (anonymous):

you buy a storage rack that holds 40 cds. you have 27 cds. write an inequality that describes how many more CDs you can buy and still have no more Cds than the rack can hold. You buy 15 Cds. will they all still fit!!!!!!!!!help please

OpenStudy (anonymous):

Well you can buy 0 CDs, or up to 13. If you buy 14 you would be over. So we get if x is the number of CDs that can be bought: \[0 \le x \le 13\] Obviously 15 CDs will not fit because 27+15=42, which is more than the rack's capacity of 40.

OpenStudy (anonymous):

thanks so much